OpenStudy (anonymous):

what number must you add to complete the square? x^2+2x=13 A. 2 B. 1 C. 3 D. 4

OpenStudy (samsungfanboy):

I'll tag some cool users that can help you

OpenStudy (anonymous):

mkay thanks

OpenStudy (samsungfanboy):

@dan815 @AlexandervonHumboldt2 @Nnesha

OpenStudy (samsungfanboy):

You're welcome! c;

Nnesha (nnesha):

x^2+2x=13 balance the equation we need to add (b/2)^2 to bht sides in order to complete the square (or to create the perfect square trinomial

Nnesha (nnesha):

\[\rm Ax^2+Bx+C=0\] this is a quadratic equation where b is the coefficient of x term

Nnesha (nnesha):

both**

OpenStudy (anonymous):

i still don't understand... o.o

Nnesha (nnesha):

well these are steps for`completing the square` 1) move the constant term to the other side of the equation 2) divide all terms by leading coefficient 3) complete the square by adding (b/2)^2 both sides

OpenStudy (anonymous):

i got it. its 1

Nnesha (nnesha):

as you can see x^2+2x=13 constant term is already at the right side and leading coefficient is one now just add (b/2)^2 both side (take half of x coefficient and then square it )

OpenStudy (anonymous):

i have another question.. solve this quadratic equation by completing the square. x2+2x=20 x=1 plus or minus square root of 24 x=-1 plus or minus square root of 24 x=-1 plus or minus square root of 21 x=1 plus or minus square root of 21

Nnesha (nnesha):

same steps how would you complete the square ? balance the equation by adding (b/2)^2 both sides